OpenStudy (anonymous):

Fan and medal!! Solve using perfect square factoring patterns. x^2 – 12x + 36 = 4 A. {–8, –4} B. {–2, 10} C. {4, 8} D. {10, –2}

OpenStudy (anonymous):

x^2 – 12x + 36 = 4 (x - 6)^2 = 2^2 square root both sides x - 6 = 2 OR x - 6 = -2 x = 8 OR x = 4

OpenStudy (anonymous):

So C? :)

OpenStudy (anonymous):

@BangkokGarrett

OpenStudy (anonymous):

@Darius25 is it C?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I have one more and I'll fan and medal for this one too!

OpenStudy (anonymous):

ok were is it

OpenStudy (anonymous):

Solve using perfect square factoring patterns. y^2 + 14y + 49 = 0 A. {0, –1} B. {7} C. {7, –7} D. {–7}

OpenStudy (anonymous):

D

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

wat grade are u in